Jarod is having a problem with rabbits getting into his vegetable garden, so he decides to fence it in. The length of the garden is 7 feet more than 5 times the width. He needs 86 feet of fencing to do the job. How many feet is the length of the garden?
Answer by rapaljer(4671)

Let x = width of the garden
5x+7 = length of the garden

2W + 2L = P
2(x) + 2(5x+7) = 86
2x+10x + 14 = 86
12x + 14= 86
12x=72
x=72/12=6 feet WIDTH
LENGTH = 5x+7
LENGTH = 5*6+ 7 = 30 +7 = 37 feet

Check:
2*6 + 2*37
12 + 74 = 86 feet

For additional explanation on word problems like this, please see my website by clicking on my tutor name "Rapaljer" anywhere in algebra.com. Look for the link on my homepage called "Basic, Intermediate, and College Algebra: One Step at a Time". Select "Basic Algebra" and look for Chapter 1. There are a couple of sections on Word Problems. See also my page "MATH IN LIVING COLOR" for solutions to many of these exercises, IN COLOR!!

R^2

Dr. Robert J. Rapalje
Seminole Community College
Altamonte Springs Campus
Florida